OpenStudy (anonymous):

Jeremy is 6 feet tall and casts a shadow that is 17 feet long. At the same instant a nearby tree casts a shadow that is 85 feet long. How tall is the tree?

OpenStudy (anonymous):

id like to know how to get the answer too! (^.^)

OpenStudy (sepeario):

We use ratios for this answer.

OpenStudy (sepeario):

17/85 = 6/x 17x = 510 x = 30

OpenStudy (anonymous):

thx! (^.^)